# Anisotropic models in LQC with GBP polymerisation

### v1

@@ III Solutions
where $M$ is a positive integration constant. It is also possible to show that the product $p_{c}c$ is constant Fernando ; Fernando2 , and this result does not depend on the lapse chosen. Indeed, from (8) and (10) we have
@@ III.2 The minimal radius
As $b^{2}_{0}>\gamma^{2}\delta_{b}^{2}$, we have $\sqrt{p_{c}}_{\text{crit}1}>\sqrt{p_{c}}_{\text{crit}2}$. In this way, we will define $r_{\text{max}}\equiv 2M$ and $r_{\text{min}}\equiv 2M\gamma^{2}\delta_{b}^{2}/b_{0}^{2}$. We see that, when $\delta_{b}\rightarrow 0$, the minimum radius goes to zero. In other words, the presence of quantum effects, encoded in the $\delta_{b}$ parameter, is responsible for preventing the initial singularity.
The polymerisation parameter can be fixed by imposing the Modesto’s minimal area condition modesto , which identifies the area of the bouncing surface with the area gap of LQG,

### v2

@@ III Solutions
where $M$ is a positive integration constant. In the context of non-singular black-hole solutions, $M$ is identified with the black/white hole mass Fernando2 . In the present case, as the KS $3$-space is finite, it is natural to identify $M$ with its energy content, as will be clearer below.
It is also possible to show that the product $p_{c}c$ is constant Fernando ; Fernando2 , and this result does not depend on the lapse chosen. Indeed, from (8) and (10) we have
@@ III.2 The minimal radius
As $b^{2}_{0}>\gamma^{2}\delta_{b}^{2}$, we have $\sqrt{p_{c}}_{\text{crit}1}>\sqrt{p_{c}}_{\text{crit}2}$. In this way, we will define $r_{\text{max}}\equiv 2M$ and $r_{\text{min}}\equiv 2M\gamma^{2}\delta_{b}^{2}/b_{0}^{2}$.
We see that, in the KS case, the minimum radius is proportional to the polymerisation parameter, and it goes to zero when $\delta_{b}\rightarrow 0$. In other words, the role of quantum effects – which are responsible for preventing the initial singularity – is encoded in the polymerisation procedure. The polymerisation parameter can be fixed by imposing the Modesto’s minimal area condition modesto , which identifies the area of the bouncing surface with the area gap of Loop Quantum Gravity,
@@ IV.4 The cosmological constant
This is a good point to discuss the interpretation of the integration constant $M$, by looking at its definition in the presence of a cosmological constant, Eq. (39). In the asymptotic de Sitter limit, the second term on the right-hand side is precisely (minus) the energy stored in the cosmological constant inside a sphere of radius $r=\sqrt{p_{c}}$, since the energy density is given by $\Lambda/(8\pi)$ and the de Sitter volume is $4\pi r^{3}/3$. Therefore, from (39) $M$ can be interpreted as the energy content in the KS space when the cosmological constant is negligible.
@@ V.1 Minimal radius
In this way, the interpretation of the integration constant $M$ is completely different in the Bianchi III context as compared to the Kantowski-Sachs case. On one hand, as the hyperbolic space is infinite, it cannot represent a total energy content. On the other hand, the minimum radius in the Bianchi III space-time originates from the postulation of the LQG minimal area, independently of any polymerisation scheme. The quantum gravity effects are, in this case, encoded only in the LQG area quantisation and, in this sense, $M$ plays the role of an effective quantum gravity parameter, in the same way as the polymerisation parameter in the Kantowski-Sachs solution. [^5] Incidentally, the value $M=\sqrt{2}/4$, found when applying the minimal area condition, coincides with the lower bound value in the KS solution, obtained from (21) in the limit $\delta_{b}\rightarrow\infty$.